# FoMo: A Foundation Model for Mobile Traffic Forecasting with Diffusion Model

### v1

@@ 4.3. Zero/Few shot learning (RQ2)
To evaluate FoMo’s zero/ few-shot learning capabilities, we select two datasets that FoMo has not encountered during training: Hangzhou (China) and Munich (Germany). We choose 4 baselines that perform well in previous multitask forecasting: KG-Diff, TimeGPT, Lagllama, and UniST. The results are shown in Figure LABEL:zero_few1, where 5% few-shot and 10% few-shot represent the model training with a small amount of data (5% and 10%, respectively). It shows that FoMo exhibits good zero-shot performance, especially in the Munich dataset, where FoMo’s zero-shot performance even surpasses that of KG-Diff after small-scale training. After training with a small amount of data, all models show varying degrees of improvement. FoMo still demonstrates the best performance, indicating that FoMo can utilize the pre-trained model to quickly capture general features within unseen mobile data. We also visualize the performance in zero/few-shot scenarios, as shown in Figure LABEL:zero_few2. We select a long-term forecasting task, with the results for $\rightarrow$ 5% $\rightarrow$ 10% few-shot from left to right in the figure. It can be observed that FoMo can learn the general distribution characteristics of mobile traffic in the zero-shot phase, and after training with a small sample, the model realizes accurate traffic forecasting.
We comprehensively summarize and compare the FoMo algorithm, UniST, Lagllama, and KG-Diff, as shown in Figure 5. In the figure, we highlight the performance of our model as well as the performance of the second-best model. FoMo consistently delivers the best performance across various tasks and multiple cities. For example, in few-shot tasks, FoMo achieves performance improvements of 10.6% and 32.1% in different cities, demonstrating the model’s universality.
Figure 5. Comparison of forecasting models in Nanjing (left) and Nanchang (right) datasets. FoMo has the best performance across diverse tasks and cities.
Figure 6. Training performance vs. model scales.
4.4. Scaling performance of FoMo (RQ3)
Scaling performance reflects the relationship between model parameters, data size, and overall performance. Understanding the scaling performance of foundation models provides valuable guidance for parameter selection during model deployment, which optimizes computational and storage overheads for the entire system. We provide 5 different model settings based on various transformer layers and hidden features: FoMo with 5M parameters (5M params), 35M FoMo (35M params), 100M FoMo (100M params), 150M FoMo (150M params), and 200M FoMo (200M params).
We first examine the scaling performance between model parameters and training loss in Figure 6, where the loss is calculated via MSE (Equation 16). The results indicate that as the model size increases, the denoising network’s capability improves, leading to enhanced learning capability of mobile traffic features. For smaller models, increasing the number of parameters significantly boosts learning capacity (as seen from 5M parameters $\rightarrow$ 35M parameters in the figure). In contrast, for larger models, a marginal diminishing becomes apparent (as depicted of 100M parameters $\rightarrow$ 200M parameters). To explore these observations, we examine the relationship between model size and performance across tasks, as shown in Figure LABEL:scale2(a). The results align with Figure 6: smaller models improve rapidly with more parameters, while larger models show minimal gains due to diminishing returns. We attribute this to parameter redundancy in larger models relative to fixed training data, leading to overfitting and limiting performance improvements. To validate the model’s scaling performance, we assess how it performs with varying dataset sizes, as shown in Figure LABEL:scale2(b). The results indicate that with small datasets, models with excessive parameters may suffer performance degradation, consistent with earlier findings. However, as data volume increases, larger models utilize their extensive parameters to improve performance. In contrast, smaller models struggle to capture diverse features, leading to a performance decline.
Discussion. From these observations, we preliminarily identify a scaling law for FoMo: For mobile traffic forecasting, merely increasing model parameters doesn’t guarantee better performance. The model size should be chosen based on the amount of collected data. This insight leads us to further investigate the quantitative relationship between model scale and factors like urban size, population, and temporal granularity to optimize performance.
@@ 4.4. Scaling performance of FoMo (RQ3)
4.5. Ablation study
We conduct ablation experiments on FoMO, as shown in Table 5, with FoMo-user and FoMo-POI representing the incorporation of mobile users and POI distributions, respectively, during the fine-tuning process. It can be observed that adding these two contextual features at the fine-tuning stage enhances model performance to varying degrees. Moreover, the performance degradation of FoMo-POI is more significant, indicating that mobile users better reflect the dynamic characteristics of mobile traffic and are more critical for mobile traffic forecasting compared to POI distribution.
@@ 5. Deployment Applications
To validate FoMo’s prediction effectiveness, we deploy the model on the Jiutian platform, an AI platform developed by China Mobile, featuring functions like scenario construction, network simulation, optimization strategy formulation, and performance evaluation, as shown in Figure 8. The Jiutian platform offers full-element network simulation capabilities, enabling efficient simulation of interactions between communication systems and user behavior. It also supports operators in developing customized algorithms and applications, deploying them into production environments, and performing product validation and testing with real network data. The platform has now been fully deployed within China Mobile, supporting network development across 31 provinces in China.
Figure 8. Optimization framework of Jiutian platform
Deployment. We select predefined urban layout and human mobility data. FoMo is deployed in the mobile traffic module, with its predictions feeding into the optimization selection module. We focus on 3 optimization scenarios (highlighted in yellow in the figure) and evaluate the performance via network coverage, throughput, and energy consumption. The training of FoMo is conducted on 4 NVIDIA A100 GPUs (80GB each) using PyTorch 2.0.1. Table 6 summarizes the model’s parameters and training/inference time per sample.
Table 6. Model deployment efficiency on JiuTian.
@@
KG-Diff 12 128 4.1M 0.12 min 0.029 min
UniST 12 512 28.2M 0.19 min 0.009 min
Lagllama 8 144 9.9M 0.21 min 0.011min
FoMo 12 128 4.9M 0.24 min 0.043 min
16 256 34.1M 0.32 min 0.054 min
20 768 217.7M 0.82 min 0.162 min
Optimization. Mobile traffic is predicted in three ways: using historical data, predicted by Lagllama, and FoMo. The optimization module then applies the same decision-making algorithms (e.g., reinforcement learning or genetic algorithms) by referring to the three predicted data to generate 3 optimization strategies (i.e., Hist-based, Lag-based, and FoMo-based), which are then evaluated in the platform via real-world mobile traffic and human mobility.
5.1. Energy efficient BS sleep mode control
@@
We consider the C-RAN scenario, where the BS achieves cell coverage by activating different numbers of RRUs (202, 2023). The strategy for BS sleep modes involves controlling the RRU operational status (activated or sleep) based on the network load. We model the problem from the perspectives of service quality, depreciation cost, and energy consumption. We set $N$ BSs, each of which has $M$ cells to serve at time $t$. Define the traffic load that a single RRU can serve as $c_{0}$, then for BS $n$, its Quality of Service (QoS) equals $Q(n)=\sum_{t}\sum_{m}max(L_{m,t}-x_{m,t}c_{0},0)/L_{m,t}$, where $L_{t,m}$ is the actual cell load and $x_{m,t}$ is the activated RRUs. Moreover, frequent switching of RRUs can lead to a reduction in the lifespan of BSs, and the depreciation yields $W(n)=\sum_{t}\sum_{m}|x_{m,t}-x_{m,t-1}|$. BS energy consumption equals $E(n)=\sum_{t}\sum_{m}\mathcal{P}[min(L_{m,t},x_{m,t}c_{0})]$ that is determined by the load at the RRU (Zhang et al., 2017), where $\mathcal{P}[L]=\alpha L+\beta(L/c_{0})$ is the energy consumption function. Therefore, the optimization objective yields:
@@
For the BS sleeping strategy, frequently adjusting RRUs over a short period is impractical. A more reasonable approach is to assess network dynamics over a longer period and develop a long-term adjustment strategy. Therefore, we leverage FoMo’s long-term prediction capability to estimate $L_{m,t}$ via the long-term prediction mask in (5). The optimization results are shown in Figure LABEL:app(a), it can be seen that the FoMo-based strategy closely matches the real data-based strategy across different metrics, with a 21.9% improvement in QoS and a 40.7% reduction in BS depreciation better than the other two. While energy consumption lags behind the comparison strategies, our algorithm aligns with that of the Real-based strategy, indicating that FoMo accurately predicts mobile traffic and effectively meets real user demands.
5.2. Traffic-aware resource allocation
We consider cellular networks involving a single cloud server and $N$ cells, where the server allocates $K$ PRBs to each cell (Wang et al., 2024; Zafar et al., 2022). We aim to optimize the average delay of each cell by allocating a limited number of PRBs across multiple time slots. We assume that the average channel quality $h_{t}$ of each cell remains constant at time slot $t$ and varies independently between different slots. The average rate in cell $n$ can be expressed as $r_{n}(t)=B\sum\nolimits_{k}\rho_{k,n}(t)log_{2}[1+p_{n}(t)h_{n}(t)/\sigma^{2}]$, where $B$ is the PRB bandwidth and $p_{n}(t)$ is the transmission power. $\rho_{k,n}(t)$ is a binary variable indicating PRB allocation ($\rho_{k,n}(t)=1$ represents PRB $k$ is allocated to cell $n$, and vice versa). Based on the PRB selection, the end-to-end delay is expressed as $d_{n}(t)=U_{n}(t)/r_{n}(t)$, where $U_{n}(t)$ is the overall transmission tasks. When there are too many transmission tasks for a cell, it might not be possible to complete all transmissions within a single time slot, leading to data backlog. Therefore, the backlog data at time $t+1$ for the cell can be expressed as $U_{n}(t+1)=max\{U_{n}(t)+P_{n}(t)-r_{n}(t),0\}$, where $P_{n}(t)$ represents the newly arriving transmission data. Our ultimate goal is to minimize the accumulated backlog data and end-to-end delay for the cells, which yields:
(18) $min\sum\nolimits_{t}\sum\nolimits_{n}U_{n}(t+1)+d_{n}(t),$
where $P_{n}(t)$ is provided by the FoMo. Compared to the BS sleep modes control, the PRB allocation primarily focuses on the user experience over a relatively short period. Therefore, we use the short-term prediction mask in (5) to generate network traffic for the near future. As shown in Figure LABEL:app(b), our FoMo-based strategy significantly reduces system latency and overall data backlog by 16.4% and 3.5%, respectively. This is because FoMo can more accurately characterize the cells’ future transmission tasks, allowing us to make decisions based on the predicted data that are closer to reality, and improve the system performance.
5.3. Coverage-oriented BS deployment
BS deployment requires an accurate assessment of the potential network demand (Dong et al., 2020). However, there is often insufficient historical traffic data in new regions, making it hard to make reasonable deployment strategies. In this case, FoMo can use information such as population and POIs in new regions to generate potential distributions of mobile traffic, thereby guiding BS deployment. Specifically, we select a target region and divide it into a 16×16 grid of square areas. FoMo employs the generation mask in (6) to generate mobile traffic. We input the number of residents and the POI distributions into FoMo, guiding it to infer the mobile traffic in the masked portions. We calculate the cumulative traffic for each area over one day (24 hours), and the BS deployment strategy is proportional to the traffic density of each area.
To validate the effectiveness, we assess the average coverage ratio across the entire region, defined as (user quantities served by BS) / (the total population of that region), with changes in regional population obtained via the population dynamics interface of the Jiutian platform. We implement two comparison schemes: the Resident population-based scheme and the POI-based scheme, as shown in Figure LABEL:app(c). It can be observed that during the night, when most people rest in residential areas, the Resident population-based scheme achieves better coverage. In contrast, during the day, when people predominantly gather around POIs, the POI-based strategy performs better. Although our Traffic generation-based scheme does not significantly outperform the other two strategies, it effectively makes a trade-off regarding dynamic human dynamics and regional functions.
@@ 6. Related Work
Mobile traffic forecasting. It can be broadly categorized into two types: prediction and generation. Mobile traffic prediction involves estimating future values using historical traffic data, while generation learns the underlying distribution of mobile traffic data relying on external contextual information and samples new data from this distribution. Early forecasting used statistical approaches or simulation techniques (Bothe et al., 2019; De Biasio et al., 2019), but these methods typically struggled to capture complex traffic patterns. With the rise of machine learning, many studies used AI for mobile traffic forecasting. For mobile traffic prediction, LSTM models have been used to capture long-term dependencies in traffic patterns (Dalgkitsis et al., 2018; Zhu et al., 2020; Wang et al., 2022). Some studies further incorporated spatial attributes into traffic prediction, with Li et al (Li et al., 2023) combining transformer and GCN to capture spatio-temporal correlations. Wu et al (Wu et al., 2022) combined GAN with GCN to capture spatial correlations across multiple cities. The MVSTGN model (Yao et al., 2023) divided urban spaces into multi-attribute graphs to capture mobile traffic features in low-dimensional space. For mobile traffic generation, early work used GANs to capture the overall distribution of mobile traffic (Ring et al., 2019; Lin et al., 2020). SpectraGAN (Xu et al., 2021) viewed cities as images, extracting POI and land use information via CNNs and incorporating it into traffic generation. Sun et al (Sun et al., 2022) added user usage features to a GAN network, improving the accuracy of traffic generation. Hui et al (Hui et al., 2023) built a city knowledge graph incorporating extensive semantic features into traffic generation models. KGDiff (Chai et al., 2024b) used a diffusion model to generate controlled traffic while considering traffic periodicity and spatial features.
@@
Foundation models. The models typically have a large scale of parameters and deep layers. They have been proven to exhibit powerful multitasking, zero/few-shot learning, and some reasoning abilities. Inspired by this, foundational models in various vertical domains have been developed. Yang et al (Yang et al., 2023a) and Hao et al proposed foundation models in financial and biomedical domains, respectively, aim to achieve various specialized tasks like investment, quantification, and gene expression. Zhang et al (Zhang et al., 2024) proposed utilizing an urban foundation model to address various applications, including urban navigation, urban Q&A, and urban planning. Notably, many foundation models for spatio-temporal forecasting have been proposed. Leveraging existing Large Language Models(LLMs), TEMPO (Cao et al., 2024) and Time-LLM (Jin et al., 2024) introduces a prompt mechanism in the pre-trained LLM for long-term forecasting by aligning features between mobile traffic and natural language tokens with a reprogramming approach. Some methods do not rely on existing language models but reconstruct spatio-temporal foundation models using transformer architectures. STEP (Shao et al., 2022) employed a pre-training model and graph neural networks to capture spatio-temporal features. LagLLama (Rasul et al., 2024) used lag indices to annotate multi-dimensional periodic features such as monthly, daily, and hourly periods. TimeGPT (Garza and Mergenthaler-Canseco, 2023) replaced the Feedforward layer in the transformer with a CNN to enhance temporal correlations. UniST (Yuan et al., 2024a) achieved spatio-temporal prediction in urban contexts by employing a memory network.
While existing models offer advantages in forecasting mobile traffic, they are often designed for specific tasks like short-term prediction or generation. In real-world deployments, network optimization involves multiple forecasting tasks across cities, requiring frequent model switching and complex adaptations, which increases deployment costs. Foundation models are well-suited for multi-task adaptation and transfer in wireless networks. A promising approach is to develop a foundation model for mobile traffic forecasting. However, no such foundation model has emerged. The main challenge lies in mastering various forecasting tasks while integrating contextual features such as human dynamics and geographical traits. This integration is essential for building a robust model that captures the intrinsic correlations between the environment, users, and traffic.

### v3

@@ 4.3. Zero/Few shot learning (RQ2)
To evaluate UoMo’s zero/ few-shot learning capabilities, we select two datasets that UoMo has not encountered during training: Hangzhou (China) and Munich (Germany). We choose 4 baselines that perform well in previous multitask forecasting: Open-Diff, TimeGPT, Lagllama, and UniST. The results are shown in Figure LABEL:zero_few1, where 5% few-shot and 10% few-shot represent the model training with a small amount of data (5% and 10%, respectively). It shows that UoMo exhibits good zero-shot performance, especially in the Munich dataset, where UoMo’s zero-shot performance even surpasses that of Open-Diff after small-scale training. After training with a small amount of data, all models show varying degrees of improvement. UoMo still demonstrates the best performance, indicating that UoMo can utilize the pre-trained model to quickly capture general features within unseen mobile data. We visualize the performance in zero/few-shot scenarios, as shown in Figure LABEL:zero_few2. We select a long-term forecasting task, with the results for $\rightarrow$ 5% $\rightarrow$ 10% few-shot from left to right in the figure. It can be observed that UoMo can learn the general distribution characteristics of mobile traffic in the zero-shot phase, and after training with a small sample, the model realizes accurate traffic forecasting. We additionally select two datasets, Fuyang (FY) and Nanning (NN), to evaluate the model. These datasets represent two Chinese cities of different scales, for which we collect mobile traffic data at an hourly granularity. As shown in Table 4, compared to the baseline method, UoMo is also capable of quickly learning the traffic patterns of new cities with limited training data, demonstrating its transferability.
4.4. Ablation study
To test the effectiveness of our proposed fine-tuning module, we conduct ablation experiments on UoMo, as shown in Table 5, with UoMo-user and UoMo-POI representing the incorporation of mobile users and POI distributions, respectively, during the fine-tuning process. It can be observed that adding these two contextual features during fine-tuning enhances model performance to varying degrees. Moreover, the performance degradation of UoMo-POI is more significant, indicating that mobile users better reflect the dynamic characteristics of mobile traffic and are more critical for mobile traffic forecasting compared to POI distribution.
@@ 4.4. Ablation study
4.5. Scaling performance of UoMo
Scaling performance reflects the relationship between model parameters, data size, and overall performance. Understanding the scaling performance of foundation models provides valuable guidance for parameter selection during model deployment, which optimizes computational and storage overheads for the entire system. We explore the relationship between model size and performance across tasks, as shown in Figure LABEL:scale2(a). Smaller models improve quickly with more parameters, while larger models show diminishing returns. We attribute this to parameter redundancy in larger models relative to fixed training data, leading to overfitting and limiting performance improvements. To assess scaling performance, we evaluate the model with varying dataset sizes in Figure LABEL:scale2(b). Larger models degrade with small datasets, but as data volume increases, they leverage their extensive parameters to improve performance. In contrast, smaller models struggle with diverse features, leading to performance drops. From these observations, we identify a scaling regulation for UoMo: Simply increasing model parameters does not guarantee better performance in mobile traffic forecasting. The optimal model size depends on the available data, prompting further investigation into the relationship between model scale and factors like urban size, population, and temporal granularity to enhance performance.
@@ 5. Deployment Applications
Deployment. To validate UoMo’s prediction effectiveness, we deploy the model on the Jiutian platform, an AI platform developed by China Mobile, featuring functions like scenario construction, network simulation, optimization strategy formulation, and performance evaluation. The Jiutian platform offers full-element network simulation capabilities, enabling efficient simulation of interactions between communication systems and user behavior. It also supports operators in developing customized algorithms and applications, deploying them into production environments, and performing product validation and testing with real network data. The platform has now been fully deployed within China Mobile, supporting network development across 31 provinces in China. We select predefined urban layout and human mobility data. UoMo is deployed in the mobile traffic module, with its predictions feeding into the optimization selection module. We focus on 3 optimization scenarios (highlighted in yellow in the figure) and evaluate the performance via network coverage, throughput, and energy consumption. The training of UoMo is conducted on 4 NVIDIA A100 GPUs (80GB each) using PyTorch 2.0.1. Table 6 summarizes the model’s parameters and training/inference time per sample.
Table 6. Model deployment efficiency on JiuTian. The training/inference time refers to the unit time per sample, which is obtained by dividing the time taken to generate a batch of data by the number of samples in the batch.
@@
Open-Diff 6 128 10M 0.09 min 0.021 min
UniST 12 512 30M 0.19 min 0.009 min
Lagllama 8 144 10M 0.21 min 0.011min
UoMo 12 128 5M 0.24 min 0.043 min
16 256 35M 0.32 min 0.054 min
20 768 200M 0.82 min 0.162 min
Optimization. Our optimization method on the Jiutian platform is shown in figure 9. First, we use UoMo or other mobile traffic forecasting algorithms to generate traffic data. The platform then uses this generated data to formulate and solve the problem of network optimization and planning. After obtaining the optimal network strategy, we input real live network traffic data into the platform to validate the optimization strategy and assess its performance.
Figure 9. Optimization workflow via UoMo.
5.1. BS deployment
We examine a BS deployment planning at the grid level within a discrete-time $\{1,...t\}$ framework. The target area consists of $N$ grids, and $M$ BSs need to be deployed across the grids. Each grid experiences varying mobile users at different times $t$, and each BS has a maximum capacity of $C$ users at any given time. The deployment strategy’s objective is to optimize the serving user count in each grid while minimizing operation cost (deploying fewer BSs in grids with fewer mobile users) and reducing capacity shortfalls (preventing situations where grids lack BSs and cannot accommodate excess users). Since there is no historical data on mobile user distribution for the target area, we rely on forecasting methods to estimate future mobile traffic demands. Three different estimation methods are utilized: the mobile traffic generation approach (UoMo-based) using the generation mask in (2), POI distribution (POI-based), and residential population distribution (Resident-based). Our focus is on addressing the BS deployment problem with a long-term perspective. The decision variables include $x_{i}$, which denotes the number of BSs deployed in grid $i$, and $u^{t}_{i}$, which represents the mobile traffic served by grid $i$ at time $t$. Then the BS deployment problem yields
(12) $max\sum_{t}^{T}\sum_{i}^{N}\Big{(}y^{t}_{i}-\beta(\hat{U}^{t}_{i}-y^{t}_{i})^{+}-\alpha\sum_{i}^{M}x_{i}\Big{)}$
[…]
where $(y)^{+}=max(0,y)$ denote choosing the maximize value between $y$ and 0, and $\hat{U}^{t}_{i}$ is the mobile network demand that is estimated by UoMo, POI, and residential distribution. We employ the Pulp library to solve the optimization problem described above, and we can derive 3 distinct BS deployment strategies by referring to the three estimation methods. Subsequently, we test the optimization strategies using real user mobility trajectory data in Nanchang, streamed by the Jiutian platform, with the test results presented in Figure LABEL:app(a). The UoMo-based strategy maximizes service revenue while significantly lowering operational costs and capacity deficits, increasing the served user ratio by 25.3% (0.652 to 0.817), and reducing costs and deficits by 18.03% (0.549 to 0.45) and 9.00% (0.733 to 0.667), respectively. This success is attributed to the UoMo takes into account the time-varying nature of network usage across different regions. By accurately estimating the mobile traffic predictions with UoMo, we can effectively capture the dynamic patterns of human activities over time in a region. In contrast, the POI-based and resident-based strategies rely solely on static attributes to guide BS deployment. These methods are less effective in capturing the dynamic activity patterns of humans over time, and fall short when compared to those based on traffic.
5.2. BS sleep control
@@
We consider the C-RAN scenario, where the BS achieves cell coverage by activating different numbers of RRUs (2023ArtificialIF, 1). BS sleep strategy involves controlling the RRU operational status (activated or sleep) based on the network load. We model the problem from the perspectives of service quality, depreciation cost, and energy consumption. We set $N$ BSs, each of which has $M$ cells to serve at time $t$. Define the traffic load that a single RRU can serve as $c_{0}$, then for BS $n$, its Quality of Service (QoS) equals $Q(n)=\sum_{t}\sum_{m}max(L_{m,t}-x_{m,t}c_{0},0)/L_{m,t}$, where $L_{t,m}$ is the actual cell load and $x_{m,t}$ is the activated RRUs. Moreover, frequent switching of RRUs can lead to a reduction in the lifespan of BSs, and the depreciation yields $W(n)=\sum_{t}\sum_{m}|x_{m,t}-x_{m,t-1}|$. BS energy consumption equals $E(n)=\sum_{t}\sum_{m}\mathcal{P}[min(L_{m,t},x_{m,t}c_{0})]$ that is determined by the load at the RRU (10.1145/3139958.3140053, 70), where $\mathcal{P}[L]=\alpha L+\beta(L/c_{0})$ is the energy consumption function. Therefore, the optimization objective yields:
@@
For the BS sleeping strategy, frequently adjusting RRUs over a short period is impractical. A more reasonable approach is to assess network dynamics over a longer period and develop a long-term adjustment strategy. Therefore, we leverage UoMo’s long-term prediction capability to estimate $L_{m,t}$ via the long-term prediction mask in (1). As shown in Figure LABEL:app(b), the UoMo-based strategy closely aligns with the Real-based one, achieving a 21.9% QoS improvement (0.645 to 0.504) and up to 40.7% lower BS depreciation (0.83 to 0.59) than the others. Although energy consumption is higher than the baselines, it matches the Real-based strategy, demonstrating UoMo’s accurate traffic prediction and strong demand alignment.
@@ 6. Related Work
Mobile traffic forecasting. It can be broadly categorized into two types: prediction and generation. Mobile traffic prediction involves estimating future values using historical data, while generation learns the underlying distribution of mobile traffic relying on external contextual information and samples new data from this distribution. Early forecasting used statistical approaches or simulation techniques (8677270, 3, 11), but these methods typically struggled to capture complex traffic patterns. With the rise of machine learning, many studies used AI for mobile traffic forecasting. For mobile traffic prediction, LSTM models have been used to capture long-term dependencies in traffic patterns (10.1145/3291533.3291540, 10, 77, 50). Some studies incorporated spatial attributes into traffic prediction, with Li et al (10.1145/3586164, 27) combining transformer and GCN to capture spatio-temporal correlations. Wu et al (9662277, 55) combined GAN with GCN to capture spatial correlations across multiple cities. The MVSTGN model (9625773, 64) divided urban spaces into multi-attribute graphs to capture mobile traffic features in latent space. For mobile traffic generation, early work used GANs to capture the overall distribution of mobile traffic (Ring_2019, 44, 30). SpectraGAN (10.1145/3485983.3494844, 58) viewed cities as images, extracting POI and land use information via CNNs and incorporating it into traffic generation. Sun et al (9858874, 46) added user usage features to a GAN network, improving the accuracy of traffic generation. Hui et al (10.1145/3580305.3599853, 21) built a city knowledge graph incorporating extensive semantic features into traffic generation models. Open-Diff (10.1145/3637528.3671544, 8) used a diffusion model to generate grid-level traffic via open contextual data.
@@
Universal and foundation models. These models show to excel in multitasking and zero/few-shot learning, and have been applied across various specialized domains. Yang et al (yang2023fingpt, 61) and Zhang et al (Zhang2024TowardsUG, 72) proposed foundation models aim to achieve various specialized tasks like investment, quantification, and urban navigation. Notably, many universal models for spatio-temporal forecasting have been proposed. Leveraging existing LLM, TEMPO (cao2024tempopromtrained, 7) and Time-LLM (jin2024timellmtimeseriesforecasting, 23) introduces a prompt mechanism in the pre-trained LLM for long-term forecasting by aligning features between mobile traffic and natural language tokens with a reprogramming approach. Some methods do not rely on existing language models but reconstruct spatio-temporal foundation models using transformer architectures. LagLLama (rasul2024lagllamafoundationmodelsprobabilistic, 43) used lag indices to annotate multi-dimensional periodic features such as monthly, daily, and hourly periods. TimeGPT (garza2023timegpt1, 15) replaced the Feedforward layer in the transformer with a CNN to enhance temporal correlations. UniST (10.1145/3637528.3671662, 69) achieved spatio-temporal prediction in urban contexts by employing a memory network. While existing models offer advantages in forecasting mobile traffic, they are often designed for specific tasks like short-term prediction or generation (ijcai2024p0921, 73, 22, 74). In real-world deployments, network optimization involves multiple forecasting tasks across cities, requiring frequent model switching and complex adaptations, which increases deployment costs. The main challenge lies in mastering various forecasting tasks while integrating contextual features such as human dynamics and geographical traits. This integration is essential for building a robust model that captures the intrinsic correlations between the environment, users, and mobile traffic.